# Data-free Asymptotics-Informed Operator Networks for Singularly Perturbed PDEs

### v1

@@ Abstract
Recent advances in machine learning (ML) have opened new possibilities for solving partial differential equations (PDEs), yet robust performance in challenging regimes remains limited. In particular, singularly perturbed differential equations exhibit sharp boundary or interior layers with rapid transitions, where standard ML surrogates often fail without extensive resolution. Generating training data for such problems is also costly, as accurate reference solutions typically require massive adaptive mesh refinement. In this work, we propose eFEONet, an enriched Finite Element Operator Network tailored to singularly perturbed problems. Guided by classical singular perturbation theory, eFEONet augments the operator-learning framework with specialized enrichment basis functions that encode the asymptotic structure of layer solutions. This design enables accurate approximation of sharp transitions without relying on large datasets, and can operate with minimal supervision—or even in a data-free manner under appropriate settings. We further provide a rigorous convergence analysis of the proposed method and demonstrate its effectiveness through extensive experiments on representative problems featuring both boundary and interior layers.
@@ 1 Introduction
The use of machine learning (ML) to solve partial differential equations (PDEs) has made significant advancements in recent years, offering innovative approaches to tackle longstanding challenges in scientific computing (lagaris1998artificial; PINN001; yu2018deep; ainsworth2021galerkin). Among these methods, operator networks have emerged as a practical and efficient tool due to their ability to infer solutions quickly after training (lu2021learning; li2021fourier). Unlike classical numerical methods, which solve PDEs iteratively for each new set of conditions, operator networks enable rapid predictions by learning the solution operator itself. This capability has made operator networks a promising new paradigm in the study of parametric PDEs. However, there remain notable limitations in applying operator networks to real-world problems. One key challenge is that training an operator network typically requires a pre-generated dataset of solutions, which are often constructed using conventional numerical methods. This process can be computationally expensive, particularly for complex PDEs. Singularly perturbed differential equations, in particular, present unique difficulties. These equations often exhibit rapid transitions within thin regions known as boundary or interior layers (see Figure 1), where constructing accurate datasets is both costly and technically challenging. Moreover, the sharp transitions inherent in thin layers can degrade the performance of operator networks, which tend to rely on smooth priors (lu2022comprehensive).
@@
Boundary and interior layer phenomena are of paramount importance in many scientific and engineering disciplines, including fluid dynamics, biology, and chemical reactions (schlichting2016boundary; batchelor2000introduction). These problems are characterized by sharp changes in solution profiles within thin layers, making them notoriously difficult to handle even with advanced numerical methods. The challenge arises from the small diffusive parameter $\varepsilon>0$ in these equations, which leads to steep gradients over small spatial regions. See Figure 2 where the examples of boundary and interior layer phenomena are presented. Developing methods to accurately and efficiently solve such problems remains a challenging task in scientific computing (FEM1; FEM2). Machine learning-based approaches face additional challenges because they are inherently better at learning smooth functions but struggle to accurately capture sharp transitions. Neural networks, for instance, are often designed to approximate solutions that vary gradually, making it difficult to capture the steep gradients and sharp transitions characteristic of boundary or interior layers (PINN007). These boundary and interior layer problems require computationally expensive massive mesh refinement to obtain accurate solutions. Moreover, as $\varepsilon>0$ decreases, the mesh size must become finer, following an approximate scaling of $\text{\it adaptive mesh size}\simeq\varepsilon$. This results in a significant drawback, as data generation can become prohibitively expensive in many cases. This limitation highlights the need for new architectures and methodologies that can handle these complexities without compromising accuracy or efficiency.
@@
Operator learning trains models to approximate PDE solution operators using datasets of input-output pairs from numerical solvers (bhatnagar2019prediction; guo2016convolutional; khoo2017solving; zhu2018bayesian), enabling efficient and real-time predictions for varying inputs (NEURIPS2020_4b21cf96). Notable architectures include the Fourier Neural Operator (FNO) (kovachki2021neural) and DeepONet (lu2021learning). Recent advances also explore message-passing frameworks to accommodate complex problem structures (brandstetter2022message; lienen2022learning; pfaff2021learning; boussif2022magnet). In addition, transformer-based architectures have been introduced (cao2021choose; wang2025cvit; pmlr-v235-hao24d), along with emerging foundation models tailored for PDEs(herde2024poseidon; ye2024pdeformer). Despite these developments, operator learning still faces challenges in generalization, data efficiency, and resolving sharp solution features. Among various operator-learning models, unsupervised physics-based operator networks incorporate governing equations directly into neural operator architectures, minimizing or completely removing the need for labeled training data. Variational frameworks such as FEONet (lee2024finite) and SCLON (SCLON) use PDE residuals in weak form to achieve accurate predictions without explicit simulation data. Similarly, physics-informed neural operator approaches like PINNs (PINN001; han2018solving), PINO (PINO), and PIDeepONet (PIDON) can also be formulated to rely entirely on PDE constraints and boundary conditions. Despite recent progress, accurately capturing multiscale phenomena and sharp gradients without labeled data remains challenging, highlighting the need for more robust unsupervised approaches.
@@
On the other hand, deep learning has emerged as a promising approach for solving singularly perturbed PDEs, with physics-informed methods also contributing to this effort (SP01; SP02). However, these approaches often lack scalability and remain effective only in limited scenarios. The study on stiff chemical kinetics (GOSWAMI2024116674) utilizes deep neural operators specifically for reaction-diffusion stiffness, limiting its applicability compared to our method, which addresses a broader class of singularly perturbed PDEs, including boundary and interior layers, particularly in data-scarce scenarios. Recently, a homotopy-based approach to learn the singularly perturbed problems was proposed by (chen2025learn) for specific PDE instances rather than operator learning approaches. ComFNO (SP03) incorporates asymptotic expansions to better handle singular perturbations. Nonetheless, challenges persist, including the need for large training datasets, difficulty in accurately capturing sharp transitions, and a lack of rigorous theoretical foundations to ensure broader reliability.
@@
In this paper, we propose an enriched Finite Element Operator Network (eFEONet), specifically designed to address these challenges. eFEONet builds upon the FEONet framework (lee2024finite), a highly data-efficient operator learning method that requires minimal training data or no dataset at all. Unlike traditional operator networks, eFEONet leverages the structure of finite element methods (FEMs), where the solution is expressed as a linear combination of nodal coefficients and basis functions. This design not only eliminates the need for large datasets but also ensures the exact satisfaction of boundary conditions. By incorporating insights from singular perturbation analysis in PDE theory, we design special basis functions within the finite element framework that capture the asymptotic beheavior of solutions in boundary or interior layers (BLA_3). This approach enables accurate modeling of sharp transitions while maintaining computational efficiency. Recently, Component Fourier Neural Operator (ComFNO) (SP03), a modified version of FNO, attempted to incorporate asymptotic behavior into its methodology. However, these approaches did not rigorously consider singular perturbation analysis, leading to suboptimal accuracy in capturing sharp layer structures. Moreover, as we can see in Section 3, despite utilizing 900 training samples, ComFNO exhibited an error that was two orders of magnitude higher than our eFEONet, which required no training data at all. We validate our approach through theoretical convergence analysis and empirical results on various singularly perturbed problems, including both boundary and interior layers (H2020; BLA_3). The results demonstrate that eFEONet achieves high accuracy and efficiency, even for problems with strong boundary or interior layer phenomena such as convection-dominated PDEs (Stynes_2005).
@@ 2 Proposed Method: Enriched Finite Element Operator Network
$-\varepsilon\,{\rm{div}}\,(\boldsymbol{a}(\mathbf{x})\nabla u_{\varepsilon})+\boldsymbol{b}(\mathbf{x})\cdot\nabla u_{\varepsilon}$ $=f\quad{\rm{in}}\,\,D.$ (1)
@@
Here we assume that the singular perturbation parameter $\varepsilon>0$ is very small so that the boundary layer phenomenon occurs. Furthermore, to highlight that the shape of a solution depends on $\varepsilon>0$, we will denote the solution as $u_{\varepsilon}$. As we will explain in more detail later, we propose an operator-learning approach for the singular perturbation problem that enables real-time solution predictions whenever the input data of the PDE varies. As a prototype model, we set the external force $f$ as an input of neural networks, and train the model so that the neural networks can learn the operator $\mathcal{G}:f\mapsto u_{\varepsilon}$. Note, however, that our method can be easily extended to various forms of input functions, including boundary conditions, variable coefficients, or initial conditions (see, e.g., (lee2024finite)).
2.1 Finite Element Method
The finite element method (FEM) is a general technique for the numerical solution of PDEs. It is based on the variational formulation of the PDE (1), which seeks to find a function $u_{\varepsilon}\in V$ satisfying
@@ 2.1 Finite Element Method
where $V$ is typically an infinite-dimensional function space for the solution and test functions. The first step in finite element method (FEM) theory is to discretize the domain $D\subset\mathbb{R}^{d}$, known as a triangulation. For $d=1$ and $D=[a,b]$, this involves points $a=x_{0}<x_{1}<\cdots<x_{K}=b,$ with each interval $x_{i-1},x_{i}]$ forming a 1-simplex. For $d=2$, the triangulation consists of closed triangles $T_{i}$ (2-simplexes), $i=1,\ldots,K$, whose interiors are disjoint. If $i\neq j$ and $T_{i}\cap T_{j}\neq\emptyset$, then the intersection is either a shared vertex or edge. For $d\geq 3$, elements are $d$-simplexes. Let $h_{T}$ denote the longest edge of a triangle $T$, and define the global mesh size as $h=\max_{T}h_{T}$. Let $S_{h}$ be the space of continuous functions $v_{h}$ on $D$ such that the restriction of $v_{h}$ to each element is a polynomial. The finite-dimensional ansatz space is then defined as $V_{h}=S_{h}\cap V$. Let $\{\mathbf{x}_{i}\}$ denote the triangulation vertices, and $\{\phi_{j}\}$ the nodal basis for $V_{h}$, where $\phi_{j}(\mathbf{x}_{i})=\delta_{ij}$. Using piecewise linear basis functions defines the P1-element method; using piecewise quadratic polynomials gives the P2-element method. The dimension of $V_{h}$ depends on the triangulation and hence on the mesh parameter $h$.
@@ 2.2 Enriched FEONet with a Corrector Basis
In this section, we shall introduce our main method, the enriched FEONet (eFEONet). One key novelty of the eFEONet is to utilize extra basis functions derived from theoretical arguments in (BLA_3). For a clear illustration of the proposed method, we shall explain it through a simple example of the following form:
@@
where $0<\varepsilon\ll 1$. As we can see from Figure 2, when $\varepsilon>0$ is small, it is difficult to expect other well-known methods including classical FEM and physics-informed neural network (PINN) to achieve good performance due to the sharp transitions near the boundary. To accurately capture the boundary layer, we incorporate an additional basis function, commonly referred to as the corrector function in mathematical analysis, for example in this case, defined as: $\phi_{\rm cor}(x):=e^{-(1+x)/\varepsilon}-(1-(1-e^{-2/\varepsilon})(x+1)/2)$. Such a basis function reflects the boundary layer properties of the given equation and is derived from theoretical arguments. The derivation of various corrector basis functions will be addressed in (BLA_3). The corrector basis is added to the standard nodal basis functions of FEM to construct an enriched Galerkin space. In other words, for enriched FEONet for the singularly-perturbed problems, we now replace the original ansatz space $V_{h}$ by the enriched Galerkin space $\overline{V}_{h}=\{\phi_{\rm cor},\phi_{1},\phi_{2},\cdots,\phi_{N(h)}\}$, where the corrector basis $\phi_{\rm cor}$ has been added to $V_{h}$. It is noteworthy that no significant additional computational cost occurs, as the enriched basis is only restricted to boundary elements. In general, neural networks assume a smooth prior, which makes them less effective in handling boundary layers. This can lead to unstable training due to the direct calculation of the PDE residual. In contrast, the eFEONet leverages theory-guided basis functions, allowing its predicted solution to precisely capture the sharp transitions near the boundary. Encapsulating the above discussion, the enriched FEM for the boundary layer problem can be written as follows: we seek $u^{\rm en}_{\varepsilon,h}\in\overline{V}_{h}$ satisfying
@@
Our framework employs corrector functions tailored to specific problem classes, yet they are not confined to individual instances. For families of PDEs with analogous singular behavior, the same correctors can often be applied effectively. In convection–diffusion equations, for example, the boundary layer typically has a thickness proportional to $\varepsilon$ with an exponential profile, a structure preserved even with additional reaction terms.
@@
Some preliminary results show that one could attempt to learn the corrector bases using data. In contrast, our approach constructs them via classical numerical analysis, which not only requires no data but also achieves substantially better performance. This integration of analytical methods into an operator learning framework constitutes the main novelty of our work, highlighting how analytic knowledge can maximize the efficiency of operator learning.

### v2

@@ Abstract
Recent advances in machine learning (ML) have opened new possibilities for solving partial differential equations (PDEs), yet robust performance in challenging regimes remains limited. In particular, singularly perturbed differential equations exhibit sharp boundary or interior layers with rapid transitions, where standard ML surrogates often fail without extensive resolution. Generating training data for such problems is also costly, as accurate reference solutions typically require massive adaptive mesh refinement. In this work, we propose eFEONet, an enriched Finite Element Operator Network tailored to singularly perturbed problems. Guided by classical singular perturbation theory, eFEONet augments the operator-learning framework with specialized enrichment basis functions that encode the asymptotic structure of layer solutions. This design enables accurate approximation of sharp transitions without relying on large datasets, and can operate with minimal supervision, or even in a data-free manner under appropriate settings. We further provide a rigorous convergence analysis of the proposed method and demonstrate its effectiveness through extensive experiments on representative problems featuring both boundary and interior layers.
@@ 1 Introduction
The use of machine learning (ML) to solve partial differential equations (PDEs) has made significant advancements in recent years, offering innovative approaches to tackle longstanding challenges in scientific computing (Lagaris et al., 1998; Lu et al., 2021b; Yu and others, 2018; Ainsworth and Dong, 2021). Among these methods, operator networks have emerged as a practical and efficient tool due to their ability to infer solutions quickly after training (Lu et al., 2021a; Li et al., 2021a). Unlike classical numerical methods, which solve PDEs iteratively for each new set of conditions, operator networks enable rapid predictions by learning the solution operator itself. This capability has made operator networks a promising new paradigm in the study of parametric PDEs. However, there remain notable limitations in applying operator networks to real-world problems. One key challenge is that training an operator network typically requires a pre-generated dataset of solutions, which are often constructed using conventional numerical methods. This process can be computationally expensive, particularly for complex PDEs. Singularly perturbed differential equations, in particular, present unique difficulties. These equations often exhibit rapid transitions within thin regions known as boundary or interior layers (see Figure 1), where constructing accurate datasets is both costly and technically challenging. Moreover, the sharp transitions inherent in thin layers can degrade the performance of operator networks, which tend to rely on smooth priors (Lu et al., 2022).
@@
Boundary and interior layer phenomena are of paramount importance in many scientific and engineering disciplines, including fluid dynamics, biology, and chemical reactions (Schlichting and Gersten, 2016; Batchelor, 2000). These problems are characterized by sharp changes in solution profiles within thin layers, making them notoriously difficult to handle even with advanced numerical methods. The challenge arises from the small diffusive parameter $\varepsilon>0$ in these equations, which leads to steep gradients over small spatial regions. See Figure 2 where the examples of boundary and interior layer phenomena are presented. Developing methods to accurately and efficiently solve such problems remains a challenging task in scientific computing (Zienkiewicz and Taylor, 2000; Hughes, 2000). Machine learning-based approaches face additional challenges because they are inherently better at learning smooth functions but struggle to accurately capture sharp transitions. Neural networks, for instance, are often designed to approximate solutions that vary gradually, making it difficult to capture the steep gradients and sharp transitions characteristic of boundary or interior layers (Karniadakis et al., 2021). These boundary and interior layer problems require computationally expensive massive mesh refinement to obtain accurate solutions. Moreover, as $\varepsilon>0$ decreases, the mesh size must become finer, following an approximate scaling of adaptive mesh size $\simeq\varepsilon$. This results in a significant drawback, as data generation can become prohibitively expensive in many cases. This limitation highlights the need for new architectures and methodologies that can handle these complexities without compromising accuracy or efficiency.
@@
Operator learning trains models to approximate PDE solution operators using datasets of input-output pairs from numerical solvers (Bhatnagar et al., 2019; Guo et al., 2016; Khoo et al., 2017; Zhu and Zabaras, 2018), enabling efficient and real-time predictions for varying inputs (Li et al., 2020). Notable architectures include the Fourier Neural Operator (FNO) (Kovachki et al., 2021) and DeepONet (Lu et al., 2021a). Recent advances also explore message-passing frameworks to accommodate complex problem structures (Brandstetter et al., 2022; Lienen and Günnemann, 2022; Pfaff et al., 2021; Boussif et al., 2022). In addition, transformer-based architectures have been introduced (Cao, 2021; Wang et al., 2025; Hao et al., 2024), along with emerging foundation models tailored for PDEs(Herde et al., 2024; Ye et al., 2024). Despite these developments, operator learning still faces challenges in generalization, data efficiency, and resolving sharp solution features. Among various operator-learning models, unsupervised physics-based operator networks incorporate governing equations directly into neural operator architectures, minimizing or completely removing the need for labeled training data. Variational frameworks such as Finite Element Operator Network (FEONet) (Lee et al., 2025) and Spectral Coefficient Learning via Operator Network (SCLON) (Choi et al., 2024) use PDE residuals in weak form to achieve accurate predictions without explicit simulation data. Similarly, physics-informed approaches like Physics-Informed Neural Networks (PINNs) (Lu et al., 2021b; Han et al., 2018), PINO (Li et al., 2021b), and PIDeepONet (Wang et al., 2021) can also be formulated to rely entirely on PDE constraints and boundary conditions. Despite recent progress, accurately capturing multiscale phenomena and sharp gradients without labeled data remains challenging, highlighting the need for more robust unsupervised approaches.
@@
In recent years, deep learning has emerged as a promising approach for solving singularly perturbed PDEs, with physics-informed methods also contributing to this line of research (Arzani et al., 2023; Tawfiq and Al-Abrahemee, 2014). However, many of these approaches suffer from limited scalability and tend to be effective only in restricted settings. For instance, the study on stiff chemical kinetics in (Goswami et al., 2024) employs deep neural operators tailored to reaction-diffusion stiffness, which narrows its applicability. In contrast, our method targets a broader class of singularly perturbed PDEs, including problems with boundary and interior layers, and is particularly suited to data-scarce regimes. More recently, (CHEN et al., 2025) proposed a homotopy-based learning strategy for singularly perturbed problems, but it focuses on specific PDE instances rather than an operator-learning framework. Component Fourier Neural Operator (ComFNO) (Li et al., 2024), a modified variant of FNO, also attempts to incorporate asymptotic behavior into the model. Nonetheless, key challenges remain, including the need for large training datasets, difficulties in accurately capturing sharp transitions, and the absence of rigorous theoretical foundations to ensure reliability across a wide range of problems.
@@
In this paper, we propose an enriched Finite Element Operator Network (eFEONet), specifically designed to address these challenges. eFEONet builds upon the FEONet framework (Lee et al., 2025), a highly data-efficient operator learning method that requires minimal training data or no dataset at all. Unlike traditional operator networks, eFEONet leverages the structure of finite element methods (FEMs), where the solution is expressed as a linear combination of nodal coefficients and basis functions. This design not only eliminates the need for large datasets but also ensures the exact satisfaction of boundary conditions. By incorporating insights from well-developed singular perturbation analysis in PDE theory, we design special basis functions within the finite element framework that capture the asymptotic behavior of solutions in boundary or interior layers (Gie et al., 2018). This approach enables accurate modeling of sharp transitions while maintaining computational efficiency. Furthermore, as demonstrated in Section 3, eFEONet significantly outperforms data-driven neural operator baselines in the singularly perturbed regime. In particular, while ComFNO is trained using 900 input-output samples, it exhibits errors that are two orders of magnitude larger than those of eFEONet, which is trained without any paired solution data. We validate our approach through theoretical convergence analysis and empirical results on various singularly perturbed problems, including both boundary and interior layers (Chekroun et al., 2020; Gie et al., 2018). The results demonstrate that eFEONet achieves high accuracy and efficiency, even for problems with strong boundary or interior layer phenomena such as convection-dominated PDEs (Stynes, 2005).
@@ 2 Proposed Method: Enriched Finite Element Operator Network
$-\varepsilon\,{\rm{div}}\,(\boldsymbol{a}(\mathbf{x})\nabla u_{\varepsilon})+\boldsymbol{b}(\mathbf{x})\cdot\nabla u_{\varepsilon}$ $=f\quad{\rm{in}}\,\,D,$ (1)
@@
Here we assume that the singular perturbation parameter $\varepsilon>0$ is very small so that the boundary layer phenomenon occurs. Furthermore, to highlight that the shape of a solution depends on $\varepsilon>0$, we will denote the solution as $u_{\varepsilon}$. As we will explain in more detail later, we propose an operator-learning approach for the singular perturbation problem that enables real-time solution predictions whenever the input data of the PDE varies. As a prototype model, we set the external force $f$ as an input of neural networks, and train the model so that the neural networks can learn the operator $\mathcal{G}:f\mapsto u_{\varepsilon}$. Note, however, that our method can be easily extended to various forms of input functions, including boundary conditions, variable coefficients, or initial conditions (see, e.g., (Lee et al., 2025)).
2.1 Finite Element Method (FEM)
The FEM is a general technique for the numerical solution of PDEs. It is based on the variational formulation of the PDE (1), which seeks to find a function $u_{\varepsilon}\in V$ satisfying
@@ 2.1 Finite Element Method (FEM)
where $V$ is typically an infinite-dimensional function space for the solution and test functions. The first step in FEM theory is to discretize the domain $D\subset\mathbb{R}^{d}$, known as a triangulation. For $d=1$ and $D=[a,b]$, this involves points $a=x_{0}<x_{1}<\cdots<x_{K}=b,$ with each interval $x_{i-1},x_{i}]$ forming a 1-simplex. For $d=2$, the triangulation consists of closed triangles $T_{i}$ (2-simplexes), $i=1,\ldots,K$, whose interiors are disjoint. If $i\neq j$ and $T_{i}\cap T_{j}\neq\emptyset$, then the intersection is either a shared vertex or edge. For $d\geq 3$, elements are $d$-simplexes. Let $h_{T}$ denote the longest edge of a triangle $T$, and define the global mesh size as $h=\max_{T}h_{T}$. Let $S_{h}$ be the space of continuous functions $v_{h}$ on $D$ such that the restriction of $v_{h}$ to each element is a polynomial. The finite-dimensional ansatz space is then defined as $V_{h}=S_{h}\cap V$. Let $\{\mathbf{x}_{i}\}$ denote the triangulation vertices, and $\{\phi_{j}\}$ the nodal basis for $V_{h}$ with $\phi_{j}(\mathbf{x}_{i})=\delta_{ij}$. Using piecewise linear basis functions defines the P1-element method; using piecewise quadratic polynomials gives the P2-element method. The dimension of $V_{h}$ depends on the triangulation and hence on the mesh parameter $h$.
@@ 2.2 Enriched FEONet with a Corrector Basis
In this section, we shall introduce our main method, the enriched FEONet (eFEONet). One key novelty of the eFEONet is to utilize extra basis functions derived from theoretical arguments in (Gie et al., 2018). For a clear illustration of the proposed method, we shall explain it through a simple example of the following form:
@@
where $0<\varepsilon\ll 1$. As shown in Figure 2, when $\varepsilon>0$ is small, it is difficult to expect other well-known methods, including classical FEM and PINN, to achieve good performance due to the sharp transitions near the boundary. To accurately capture the boundary layer, we incorporate an additional basis function, commonly referred to as the corrector function in mathematical analysis, for example in this case, defined as: $\phi_{\rm cor}(x):=e^{-(1+x)/\varepsilon}-(1-(1-e^{-2/\varepsilon})(x+1)/2)$. Such a basis function reflects the boundary layer properties of the given equation and is derived from theoretical arguments. The derivation of various corrector basis functions can be found in a range of references, including (Gie et al., 2018). The corrector basis is added to the standard nodal basis functions of FEM to construct an enriched Galerkin space. In other words, for enriched FEONet for the singularly-perturbed problems, we now replace the original ansatz space $V_{h}$ by the enriched Galerkin space $\overline{V}_{h}:=\{\phi_{\rm cor},\phi_{1},\phi_{2},\cdots,\phi_{N(h)}\}$, where the corrector basis $\phi_{\rm cor}$ has been added to $V_{h}$. It is noteworthy that no high additional computational cost occurs, as the enriched basis is only restricted to boundary elements. In general, neural networks assume a smooth prior, which makes them less effective in handling boundary layers. This can lead to unstable training due to the direct calculation of the PDE residual. In contrast, the eFEONet leverages theory-guided basis functions, allowing its predicted solution to capture the sharp transitions near the boundary precisely. Encapsulating the above discussion, the enriched FEM for the boundary layer problem can be written as follows: we seek $u^{\rm en}_{\varepsilon,h}\in\overline{V}_{h}$ satisfying
@@
Our framework employs corrector functions tailored to specific problem classes, yet they are not confined to individual instances. For families of PDEs with analogous singular behavior, the same correctors can often be applied effectively. In convection-diffusion equations, for example, the boundary layer typically has a thickness proportional to $\varepsilon$ with an exponential profile, a structure preserved even with additional reaction terms.
@@
Some preliminary results show that one could attempt to learn the corrector bases using data. In contrast, our approach constructs them via classical numerical analysis, which not only requires no data but also achieves substantially better performance. This integration of analytical methods into an operator-learning framework constitutes the main novelty of our work, highlighting how analytic knowledge can maximize the efficiency of operator learning.